You are ordered to give 10mcg/kg/min of dobutamine to a patient. He weighs 145 pounds. Dobutamine comes in a 12.5mg/ml package. It contains 100 ml. What is your flow rate?
Answer #1

ANSWER.

* Calculation of flow rate.

       Flow rate(ml/hr) = Volume(ml) / Time(hr).

Calculation of volume(ml).

        Volume(ml) = Desired dose/Dose in hand * Quantity.

    Here

        Ordered dose = 10 mcg/kg/min.

    Convert mcg to mg for calculation

                1 mcg = 0.001 mg

        Ordered dose = 10 * 0.001

                               = 0.01mg/kg/min.

        Weight of the patient = 145 pounds.

     Convert pound to kg for calculation

              1 pound = 0.454 kg

         Weight of this patient in kg = 145 * 0.454

                                                     = 65.8 kg.

         So

               Desired dose(D) = 0.01 * 65.8

                                           = 0.7 mg.

               Dose in hand (H) = 12.5 mg.

                     Quantity(Q) = 1 ml.

       As per above formula for volume(ml)

                Volume(ml) = D/H * Q

                                    = 0.7/12.5 * 1

                                    = 0.06 * 1

                                    = 0.06 ml.

                Time(min) = 1 min.

      Convert ml/min to ml/hr by multiply with 60.

                Rate(ml/hr) = 0.06 * 60

                                    = 3.6 ml/hr.